A typical copper might have 2*10^21 free electrons in 1cm of it's length. Suppose that the drift speed of the electrons along the wire is 0.05cm/s
a. How many electrons will pass through a cross section wire each second
b. How large a current would be flowing in the wire
1
Expert's answer
2021-07-21T09:51:43-0400

Gives

Number of free electrons (n)="2\times10^{21}"

Length(l)=1Cm

V=0.05cm/sec

Number per second

="\frac{number}{length}\times velocity"


"=\frac{2\times10^{21}}{1}\times0.05=1\times10^{20} electron/sec"

(b)

"I=\frac{Q}{t}=\frac{1\times10^{20}\times1.6\times10^{-19}}{1}=16A"